40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 13 22:32:54 2015, Koji, Summary, PSL, PMC LO degraded, usual ERA-5 replacement, LO recovered PMC_LO.pngIMG_2093.JPGIMG_2091.JPGIMG_2095.JPGIMG_2096.JPG
    Reply  Sun Nov 15 22:43:48 2015, Koji, Summary, PSL, PMC LO degraded, usual ERA-5 replacement, LO recovered C1-MULTI_E6875C_TIMESERIES-1131408017-86400.pngC1-MULTI_E6875C_TIMESERIES-1131580817-86400.png
       Reply  Mon Nov 16 19:05:59 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
          Reply  Mon Nov 16 21:32:49 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
             Reply  Tue Nov 17 16:21:10 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
                Reply  Wed Nov 18 16:51:58 2015, Koji, Summary, PSL, PMC servo calibration PMC_err_cal.pdf
                   Reply  Wed Nov 18 16:53:17 2015, Koji, Summary, PSL, PMC PZT capacitance 
                      Reply  Wed Nov 18 17:09:22 2015, Koji, Summary, PSL, PMC Servo analysis PMC_OLTF.pdfPMC_DSP.pdf
                         Reply  Fri Nov 20 15:44:12 2015, Koji, Summary, PSL, Added 17.5kHz LPF to the PMC servo PMC_OLTF.pdfPMC_noise_comparison.pdf
Message ID: 11781     Entry time: Wed Nov 18 16:53:17 2015     In reply to: 11780     Reply to this: 11782
Author: Koji 
Type: Summary 
Category: PSL 
Subject: PMC PZT capacitance 

The PMC PZT capacitance was measured.
- Turn off the HV supplies. Disconnect HV OUT cable.
- Make sure the cable is discharged.
- Measure the capacity at the cable end with an impedance meter.
=> The PMC PZT capacitance at the cable end was measured to be 222nF
Combined with the output impedance of 63.3kOhm, the LPF pole is at 11.3Hz

ELOG V3.1.3-